Can you please assist me with this problem.
Find all real solutions of the equation. (x^2+1)^2 - 18(x^2+1) = - 17.

Answer by mananth(16949)

Can you please assist me with this problem.
Find all real solutions of the equation.

(x^2+1)^2 - 18(x^2+1) = - 17. substitute x^2+1 by a
a^2-18a+17=0
a^2-17a-a+17=0
a(a-17)-1(a-17)=0
(a-17)(a-1)=0
(x^2+1-17)(x^2+1-1)
(x^2-16)*x^2
x^2(x+4)(x-4)